# Revisiting Weight Averaging for Model Merging

### v1

@@ Appendix B Further Analysis on Loss Landscapes
To directly compare the behaviors of CART and task arithmetic in the parameter space, we visualize the loss landscapes of them with respect to the multi-task loss (e.g., $\mathcal{L}=\sum_{t=1}^{T}\mathcal{L}_{t}$). As observed in the left panel of Figure 7, CART can closely approach the common basin induced by the multi-task loss, with its trajectory rotating from $\theta_{\text{avg}}$ as the rank increases. In contrast, for task arithmetic, we observed that even as the rank increases (up to full rank), it converges to a point farther from the common basin than CART, which is consistent with Figure 1. In the right panel of Figure 7, we examine the trajectory of task arithmetic with varying $\lambda$. We observe that when $\lambda=0.125$, the trajectory moves toward the weight average, and as $\lambda$ increases (e.g., to 1), there is a tendency to move away from the common basin. This trend aligns with the results presented by Ilharco et al. [12].
@@
Task-wise Loss Landscapes
In Figure 8, we investigate the loss landscape around the fine-tuned weights $\theta_{t}$ for each task. We confirm that the $\theta_{t}$ are located within the basin for most tasks. Additionally, by examining the positions of CART and task arithmetic, we observe that CART is closer to the loss basin of the single tasks, which is consistent with the results shown in Figure 5(a).

### v2

@@ Appendix E Further Analysis on Loss Landscapes
To directly compare the behaviors of CART and task arithmetic in the parameter space, we visualize the loss landscapes of them with respect to the multi-task loss (e.g., $\mathcal{L}=\sum_{t=1}^{T}\mathcal{L}_{t}$). As observed in the left panel of Figure 11, CART can closely approach the common basin induced by the multi-task loss, with its trajectory rotating from $\theta_{\text{avg}}$ as the rank increases. In contrast, for task arithmetic, we observed that even as the rank increases (up to full rank), it converges to a point farther from the common basin than CART, which is consistent with Figure 1. In the right panel of Figure 11, we examine the trajectory of task arithmetic with varying $\lambda$. We observe that when $\lambda=0.125$, the trajectory moves toward the weight average, and as $\lambda$ increases (e.g., to 1), there is a tendency to move away from the common basin.
Task-wise Loss Landscapes
In Figure 12, we investigate the loss landscape around the fine-tuned weights $\theta_{t}$ for each task. We confirm that the $\theta_{t}$ are located within the basin for most tasks. Additionally, by examining the positions of CART and task arithmetic, we observe that CART is closer to the loss basin of the single tasks, which is consistent with the results shown in Figure 10.